Please help.
Find the exact solutions for 5x^2+17x+6=0

Answer by stanbon(75887)

Find the exact solutions for 5x^2+17x+6=0
-----------------
Use the AC Method to factor:
5x^2 + 15x + 2x + 6 = 0
5x(x+3) + 2(x+3) = 0
(x+3)(5x+2) = 0
x = -3 or x = -2/5
==============
Cheers,
Stan H.